OpenStudy (anonymous):

Find all numbers for which the expression is undefined t^3-3t/t^2-81

OpenStudy (anonymous):

once again only when dividing by 0, t^2-81=0 (t-9)(t+9)=0, so when t= 9 or -9 you have a problem

OpenStudy (anonymous):

yep can't have a zero in the denominator or it's undefined

OpenStudy (anonymous):

just factor like tutor did, then set it equal to zero